CMS Corporation's balance sheet as of today is as follows:

Long-term debt (bonds, at par) $10,000,000
Preferred stock 2,000,000
Common stock ($10 par) 10,000,000
Retained earnings 4,000,000
Total debt and equity $26,000,000

The bonds have a 4.7% coupon rate, payable semiannually, and a par value of $1,000. They mature exactly 10 years from today. The yield to maturity is 12%, so the bonds now sell below par. What is the current market value of the firm's debt?

Select the correct answer.

a. $5,812,790
b. $5,814,167
c. $5,814,856
d. $5,815,545
e. $5,813,479
Answer #1

The price of one bond is:-

=PV(rate,nper,pmt,fv)

=PV(12%/2,10*2,-4.7%/2*1000,-1000)

=581.35

Total market value =581.35*10000000/1000 =$5,813,479